Math 125 Proficiency Quiz (20 minutes) Name Student number Quiz section:

1. Simplify \sqrt{8} + \sqrt{8}.

2. Evaluate

(a) sin \frac{π}{3} =

(b) cos 0 =

(c) tan \frac{π}{4} =.

3. Write down the double angle formula for sin(2x) and cos(2x).

4. Find the derivative of the following functions:

(a) ln(x^{2}):

(b) e^{x} sin x:

(c) \frac{tan x}{x+1}:

(d) sec x: